OpenStudy (anonymous):

Identify the 35th term of an arithmetic sequence where a1 = -7 and a18 = 95.

OpenStudy (anonymous):

18 is a subscript btw.

OpenStudy (amistre64):

this is sometimes simpler if we start the index at a0 a0 = -7 + d(0) a1 = -7 + d(1) a2 = -7 + d(2) a3 = -7 + d(3) a4 = -7 + d(4) etc ... our goal is to determine the value of d

OpenStudy (amistre64):

a18 then shifts its index as a17 and we have: 95 = -7 + d(17)

OpenStudy (amistre64):

our goal is to determine the value of a35 shifted is: a34 a34 = -7 + d(34)

OpenStudy (anonymous):

so when it shifts, it just goes down by one?

OpenStudy (amistre64):

you can shift it by any amount your heart desires; i chose to shift it down by 1 is all to avoid the whole d(n-1) construction in favor of dn

OpenStudy (amistre64):

spose your given a12 and a57 we would then whose to shift it down by 12 ....

OpenStudy (anonymous):

1)203 2)197 3.)168 4.) 163 these are my choices just don't know now i would get to a number with no variable

OpenStudy (amistre64):

*then chose...

OpenStudy (amistre64):

solve for d: 95 = -7 + d(17) then use it to solve for: a34 = -7 + d(34)

OpenStudy (anonymous):

can you explain further? please, sorry I'm like really bad with this

OpenStudy (amistre64):

its just algebra ... 95 = -7 + d(17) ; what value of d (or would you perfer it to be called x?) makes this work

OpenStudy (anonymous):

i got 6!

OpenStudy (amistre64):

a34 = -7 +6(34 )

OpenStudy (anonymous):

a34=197

OpenStudy (amistre64):

excellent :)

OpenStudy (anonymous):

thank you so much! i get this now!!!

OpenStudy (amistre64):

practice helps :)

OpenStudy (anonymous):

just got 100 on my quiz! whooooo